OpenStudy (kainui):

What is the property x*anything = x called? For instance infinity * 2 = infinity

OpenStudy (aravindg):

Are you referring to the identity property?

OpenStudy (mertsj):

Multiplicative Identity Property

OpenStudy (kainui):

Nearly everything is the multiplicative identity of infinity haha, alright I guess that works. I just thought it was a bit different and had some other name, thanks.

OpenStudy (samigupta8):

Hold on! Multiplicative identity is the property when we get the same number when it's multiplied with 1 and not any other number..Isn't it?? And here in your question, you remarked that x multiplied with anything gives x .So how is it going to be that identity?

OpenStudy (kainui):

I see what you're asking. Here, I'll show what I mean: \[a*b =a\] this is true when \(b=1\) for all \(a\). which is what you're thinking. But this is also true if \(a=\infty\) and for all \(b > 0\) too.

OpenStudy (samigupta8):

Correct ! But not for any other value of 'a' than infinity

OpenStudy (samigupta8):

Ain't it?

OpenStudy (kainui):

\(-\infty\) I guess haha but past that I don't know. Depends on how serious you want to get with this, if a and b were matrices we could easily end up with some other alternatives.

OpenStudy (samigupta8):

Ah! I see .Here, i were restricting myself to the numbers though!

OpenStudy (bobo-i-bo):

Hmmm, look up the "zero product property". Might be something like that.

OpenStudy (samigupta8):

Well! Still kai do you think that this property is multipliactive identity?

OpenStudy (samigupta8):

I completely agree with your notion that infinity multiplied with a number gives infinity but this is not what multiplicative identity refers to .

OpenStudy (samigupta8):

@bobo-i-bo what about that instance kainui adressed to ? I mean when infinity is multiplied with any number what will you call that property then?

OpenStudy (kainui):

Yeah I don't quite think that it's really what multiplicative identity refers to either. Beats me though, it's not too important I just wanted to know what to call this since I wanted to explain this idea to someone for a derivation with convolutions.

OpenStudy (samigupta8):

Sounds good then! Whatever it is you might have been quite successful in explaining the better understanding of this property though!

OpenStudy (bobo-i-bo):

Hmmm, not sure it completely captures the idea in question... I wish to say something more along the lines of "singular multiplication/addition". '~'